OpenStudy (anonymous):

Find three consecutive integers such that the sum of twice the smallest and 3 times the largest is 126

OpenStudy (anonymous):

Ok lets describe and label the consecutive integers so x is the smalles X+1 is the seconds consecutive integer and the third one, the largest is X+1+1=x+2

OpenStudy (anonymous):

Two times the smallest Number +3 Times the largest integer =126 Using the information i have provided above solve for x

OpenStudy (anonymous):

i still dont get it, its too hard

OpenStudy (anonymous):

do you know what consecutive integers mean?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

It means one number after the other so for example three consecutive integers are 3,4,5 or 5,6,7 So we wanna find 3 consecutive integers that when multiply the smallest number and add it to 3 times the largest number the sum will be 126 2 X Smallest Integer + 3 X Largest Integer =126 so the smallest integer is x. That x just represent any integer x+1 is the next integer and x+1+1=x+2 is the largest of the consecutive integers Therefore we have an equation of the following 2(x)+3(x+2)=126 Now solve for x

OpenStudy (anonymous):

i got 24 n thankss i understand it alot more

OpenStudy (anonymous):

ok x=24 Meaning that the smallest integer is 24 therefore the 3 consecutive integers are 24,25,26